Breeders at the American Speedy Squirrels’ Association (ASSA) breeding station in Logan, Oregon, are interested in producing “speedy” racing squirrels to compete for popularity with Iditarod in Alaska. Listed in the table below is the speed distribution of their initial population.

a. What is the mean speed of the squirrels?

b. What is the variance in the speed of the squirrels?

Squirrel speed (m/s) No. squirrels with the speed

2 1

3 4

4 7

5 5

6 2
Answer #1

a.
Mean = (1 + 4 + 7 + 5 + 2)/6
= 19/6
= 3.166 m/s

b.
To find variance,
Find the mean of data.
Subtract each number from the mean value.
Square the result.
Add the results together.
Divide the result by n-1 (n = number of data points)

= [(1 - 3.16)^2 + (4 - 3.16)^2 + (7 - 3.16)^2 + (5 - 3.16)^2 + (2 - 3.16)^2]/(6-1)
= [4.66 + 0.7056 + 14.74 + 3.38 + 1.34]/5
= 24.825/5
= 4.96